OpenStudy (anonymous):

A metal conduit will be used as a pathway for wiring through a concrete block. The conduit is a 6 foot long rod with an outer diameter of 2.8 inches and an inner diameter of 2.6 inches. What is the volume of metal in the conduit in cubic centimeters?

OpenStudy (anonymous):

You can imagine the conduit as two cylinders, the conduit as a whole and just the inside of it. If we want to get the volume of only the metal part we just get the whole volume and minus the volume of the inside (air part) leaving just the outside metal part. Use the formula V=(pi)*(r^2)*L where r is the radius and L is the length of the tube, try it yourself

OpenStudy (jhannybean):



OpenStudy (anonymous):

wow!!!

OpenStudy (johnweldon1993):

Right, so in that -lovely- diagram drawn by Jhanny up there  We need to find the Volume of the inner cylinder and subtract that from the WHOLE volume Focus on the inner cylinder first...Knowing that Volume = pi r^2 * L And also knowing radius = (1/2)diameter \(\large V_{inner} = \pi r^2 L = \pi (1.3in)^2 * 72in = ?\) And now for the whole cylinder \(\large V_{outer} = \pi r^2 L = \pi (1.4in)^2 * 72in = ?\) Finally we just need to subtract the inner volume from the whole volume to find out how much metal is there not including the giant gaping hole \[\large V_{metal} = V_{outer} - V_{inner} = ?\] **Also, once you have your answer, you need to convert it from cubic inches to cubic centimeters**